# LeTac-MPC: Learning Model Predictive Control for Tactile-reactive Grasping

### v1

@@ Abstract
Grasping is a crucial task in robotics, necessitating tactile feedback and reactive grasping adjustments for robust grasping of objects under various conditions and with differing physical properties. In this paper, we introduce LeTac-MPC, a learning-based model predictive control (MPC) for tactile-reactive grasping. Our approach enables the gripper grasp objects with different physical properties on dynamic and force-interactive tasks. We utilize a vision-based tactile sensor, GelSight [1], which is capable of perceiving high-resolution tactile feedback that contains the information of physical properties and states of the grasped object. LeTac-MPC incorporates a differentiable MPC layer designed to model the embeddings extracted by a neural network (NN) from tactile feedback. This design facilitates convergent and robust grasping control at a frequency of 25 Hz. We propose a fully automated data collection pipeline and collect a dataset only using standardized blocks with different physical properties. However, our trained controller can generalize to daily objects with different sizes, shapes, materials, and textures. Experimental results demonstrate the effectiveness and robustness of the proposed approach. We compare LeTac-MPC with two purely model-based tactile-reactive controllers (MPC and PD) and open-loop grasping. Our results show that LeTac-MPC has the best performance on dynamic and force-interactive tasks and the best generalization ability. We release our code and dataset at https://github.com/ZhengtongXu/LeTac-MPC.
@@ I Introduction
Grasping is a reactive task, in which humans estimate the states and physical properties of the grasped object by tactile feedback of fingers, and dynamically adjust finger behavior. This allows for robust grasping regardless of the object’s stiffness, whether it needs to be stationary or dynamically moved, and whether it is subjected to external forces. An ideal grasping algorithm for robots should be able to do the same. In this paper, we study the task of robotic tactile-reactive grasping. We specifically focus on a vision-based tactile sensor, GelSight [1], due to its advantages of perceiving high-resolution tactile feedback that contains rich information of physical properties and states of the grasped object.
@@
1) Complex object physical properties: Objects may have different physical properties, which poses great challenges to tactile-based perception. Existing tactile perception methods often assume that grasped objects are rigid [10, 11], and sensors, such as vision-based tactile sensors [12, 13, 14], may not be sensitive if the grasped object is softer than the sensor elastomer [1]. As a result, it can be difficult for controllers to get high-quality and stable feedback signals by model-based extraction methods when handling objects with different physical properties, and it can be challenging to generalize a same set of controller parameters to different objects since their physical properties and dynamics may different.
@@
2) Dynamic and force-interactive tasks: Statically grasping an object is often insufficient for dynamic and force-interactive tasks. For example, when robots use gripper to pick fruit, there is force interaction between the gripper, the fruit, and the branches. Similarly, in dynamic pick-and-place tasks, the grasped object may experience the inertia force and collide with unexpected obstacles during transportation. In both of these scenarios, if a gripper is not responsive enough to state changes, external disturbances, and force interactions, it can result in the grasped object falling off from the gripper. Moreover, if a larger grasping force is used for a more robust grasp, delicate objects such as fruits, eggs, bread, and crackers may be destroyed. Therefore, it is challenging to design a reactive controller to grasp with appropriate force according to the physical properties and states of the object.
@@
3) High-resolution tactile feedback integration: Another challenge is the integration of high-resolution tactile feedback, such as vision-based tactile sensor GelSight [1], which contains rich information, into a real-time control loop for tactile-reactive grasping. Traditional control methods are based on low-dimensional feedback signals, which require the extraction of required information from tactile images for dimensionality reduction [15, 16]. However, many feature extraction methods make assumptions about the physical properties and shape of the object [10, 11, 15], making it difficult to generalize to objects with different physical properties and shapes. On the other hand, robot learning methods are based on high-dimensional observations such as visual and tactile images. However, these methods often overlook crucial aspects of the control process, such as convergence, response speed, control frequency, and constraints. As a result, they are more suitable for generating open-loop or low-frequency actions, but not well-suited for tasks that demand rapid and reactive responses.
@@
To address the these challenges, we propose LeTac-MPC, a learning-based model predictive control (MPC) for tactile-reactive grasping. The novelties of the proposed approach are as follows.
@@
1) We design a neural network (NN) architecture with a differentiable MPC layer as the output layer. The MPC layer is modeled according to the control objectives to rationalize the embeddings extracted by the NN from the tactile feedback. The combination of NN and MPC layer provides the advantages of strong generalization ability and ease of implementation as a real-time controller. We implement the trained NN and MPC layer to a model predictive controller that is convergent and can run at a frequency of 25 Hz. This controller enables robust reactive grasping of various daily objects on dynamic and force-interactive tasks.
@@
2) To train the model, we use four standardized blocks of different materials with different physical properties to collect a dataset and propose a fully automated data collection pipeline. Finally, we show that despite our data collection with standardized blocks in terms of sizes, shapes, and materials, and with no textures on the block surface, our trained controller can generalize to various daily objects with different sizes, shapes, materials, and textures. This shows the strong generalization ability of LeTac-MPC.
@@
3) We also compare two purely model-based tactile-reactive controllers (MPC and PD) as baseline methods. Through experimental comparisons between these two model-based controllers, LeTac-MPC, and open-loop grasping, we show that LeTac-MPC has the best performance on dynamic and force-interactive tasks and the best generalization ability.
@@ II-A Tactile Control
Previous research has shown that tactile feedback provided by vision-based tactile sensors can be used to design manipulation controllers for rigid objects [11, 17, 16], cables [15], and clothes [18]. These methods extract state estimations of the manipulated object, such as pose and contact line with the environment, from tactile feedback and then design controllers. However, these approaches are not generalizable to objects with different physical properties due to limitations in their model-based state extraction methods. The work presented in [19] proposes a tactile control framework for robotic pushing that utilizes tactile feedback to estimate and predict state information. In [20], a learning-based method for page flipping via tactile sensing is proposed. However, these two works do not fully use the rich information about the object’s physical properties that can be obtained from tactile feedback, which is critical for enhancing generalization to different objects.
Recently, The work in [21] proposes a method to learn soft tactile sensor membrane dynamics. However, this method does not consider the dynamics of the manipulated object, which is important for deformable object manipulation. The work in [22] proposes a learning-based tactile MPC framework. However, this framework is not suitable for real-time control tasks because it uses a video prediction model, resulting in high computational cost. As mentioned in the paper, during the experimental testing, this MPC only runs at around 1 Hz.
@@ II-B Tactile Grasping
Multiple works exploit tactile feedback and NN to predict grasp success and stability [23, 24, 25, 26]. Based on the predictive ability of NN, the works in [27, 28, 29, 30] sample potential grasping actions and select the action with the lowest cost as the optimal grasping. However, these methods do not primarily focus on real-time control. They do not consider important aspects such as convergence, response speed, control frequency, and constraints in the control process. Consequently, they are better suited for low-frequency action generation rather than real-time control. For example, in [29], the computation times of different NN models for one trial are both more than 0.28 s. As a result, these methods cannot respond fast and smooth enough when the robot performs dynamic and force-interactive tasks. The works in [31, 10] maintain stable grasping by slip detection. Therefore, the accuracy of slip detection determines the performance of the method. However, in practical situations, slip detection may be insufficiently sensitive due to the different shapes and physical properties of grasped objects.
In [32, 15], grasping controllers with model-based tactile feature extraction are proposed. However, these methods cannot be widely applicable to objects with different shapes and physical properties since some feature extraction methods only work on specific types of objects. For example, GelSight, as used in [32, 15], cannot obtain obvious features for objects softer than the sensor’s gel [1]. These features are extremely difficult to process and extract using model-based methods. Instead, existing work about tactile perception demonstrates that NN can effectively extract these features [9]. Therefore, extending the learning-based tactile perception to develop a learning-based tactile control method that can generalize to objects with different shapes, textures, and physical properties is crucial.
@@ II-C Learning Physical Properties for Manipulation
Learning-based methods that incorporate tactile and visual feedback can achieve dynamic manipulation without prior knowledge of the physical properties of the manipulated object [33, 34, 35]. However, these methods are not suitable for tasks that require reactive behavior. In contrast, many manipulation tasks require real-time sensing and real-time control. In [36, 37], the authors propose methods utilizing neural radiance fields (NeRF) for manipulating objects with complex properties. However, these methods are not suitable for tasks demanding real-time control and fast response due to the relatively high computational costs. In [38], a model learning method for elastic deformable linear objects is proposed that enables real-time cable manipulation. The input of its model is the cable state extracted from the raw image, requiring markers on the cable for tracking. However, in practice, extracting cable states from images without markers is challenging. In brief, achieving real-time, learning-based robot control directly using high-dimensional images as input remains a significant challenge.
@@ III Approach
This section outlines our proposed learning-based MPC for tactile-reactive grasping. Initially, we design a NN architecture that includes a differentiable MPC layer (Sections III-A and III-B) to extract and represent the diverse and complex physical properties of grasped objects. Then, we introduce the details of training our proposed model (Section III-C). Furthermore, we describe our automated data collection pipeline in Section III-D. Finally, we present the implementation of the NN as a real-time controller in Section III-F.
By our method, the gripper can successfully grasp objects with different physical properties and shapes while applying appropriate force to avoid damage to fragile and delicate items. Our method also can adjust the grasping in real-time based on tactile feedback to improve the grasping robustness when performing dynamic and force-interactive tasks.
@@ III-A Tactile Information Encoding
We use GelSight [1], a vision-based tactile sensor that provides high-resolution images of the contact surface geometry and strain field, as shown in Fig. LABEL:fig:firstPage. To extract states and physical properties of grasped objects, we use a convolutional neural network (CNN) with a multi-layer perceptron (MLP) to encode a tactile image into a low-dimensional embedding $\mathbf{f}\in\mathbb{R}^{M}$, where $M$ is the dimension of the embedding. The works on tactile perception in [9, 40] demonstrate the effective feature extraction from tactile images using pre-trained visual models. In our implementation, we use a pre-trained ResNet-152 [41] as the CNN architecture and replace the last layer with a two-layer MLP with ReLU activation.
@@ III-B Differentiable MPC Layer
We formulate a MPC layer to use the extracted embedding for controller design, as shown in Fig. 1. In this paper, we consider the two-finger gripper with one degree of freedom linear motion. This kind of gripper is widely used due to practicality. Denoting the sampling time interval as $\Delta t$, we can first formulate the gripper motion model as:
@@
Scalars $p,v$, and $a$ are the position, velocity, and acceleration of the motion of gripper fingers. Note that $p,v$, and $a$ do not refer to the motion of the gripper itself in Cartesian space, but rather denote the one-degree-of-freedom motion of the gripper fingers. The right subscript $n$ represents the $n$-th time step.
@@
where $\mathbf{A}_{f}\in\mathbb{R}^{M}$ would be learned. This modeling approach does not lose its applicability to complex dynamics because the tactile embeddings are actually generated by a NN, which has strong nonlinear fitting capabilities. Moreover, for MPC, we do not expect the model to maintain high accuracy over long timescales because MPC operates at a high frequency with receding horizon control. Therefore, as long as the model demonstrates adequate representational capability in a short-term local area, it is sufficient for MPC performance. The combination of the NN and the MPC layer provides the advantages of strong generalization ability and ease of implementation as a real-time controller. The forward pass of the MPC layer involves solving the following optimization problem:
@@
$n$ denotes the current time step, and we can define the initial time step as $n=0$. As time progresses, there is no upper limit to $n$. Therefore, $n=0,1,2,3,\ldots$. $\mathbf{a}_{n}=[a_{n},a_{n+1},\ldots{},a_{n+N-1}]^{T}\in\mathbb{R}^{N}$ is the acceleration sequence, $N$ is the prediction horizon, the vector $\mathbf{A}_{f}\in\mathbb{R}^{M}$ is part of the state transition matrix, $\mathbf{Q}_{f}\in\mathbb{R}^{M\times M}$ is the weight matrix for $\mathbf{f}_{n}$, and $Q_{v}$ and $Q_{a}$ are the weight coefficients for $v_{n}$ and $a_{n}$, respectively. Scalar $P$ is used to amplify the terminal cost to speed up convergence. Note that the “$n$” in $v_{n},v_{n+1}$ represents the current time step $n$, while the “$N$” in $v_{n+N}$ represents the MPC horizon length $N$. Importantly, we design $\mathbf{f}_{n+1}$ is only directly related with $\mathbf{f}_{n}$ and $v_{n}$. This design helps the model avoid the influence of the different sizes of grasped objects. Therefore, the trained model can generalize to objects with different sizes naturally.

### v2

@@ Abstract
Grasping is a crucial task in robotics, necessitating tactile feedback and reactive grasping adjustments for robust grasping of objects under various conditions and with differing physical properties. In this paper, we introduce LeTac-MPC, a learning-based model predictive control (MPC) for tactile-reactive grasping. Our approach enables the gripper to grasp objects with different physical properties on dynamic and force-interactive tasks. We utilize a vision-based tactile sensor, GelSight [1], which is capable of perceiving high-resolution tactile feedback that contains information on the physical properties and states of the grasped object. LeTac-MPC incorporates a differentiable MPC layer designed to model the embeddings extracted by a neural network (NN) from tactile feedback. This design facilitates convergent and robust grasping control at a frequency of 25 Hz. We propose a fully automated data collection pipeline and collect a dataset only using standardized blocks with different physical properties. However, our trained controller can generalize to daily objects with different sizes, shapes, materials, and textures. The experimental results demonstrate the effectiveness and robustness of the proposed approach. We compare LeTac-MPC with two purely model-based tactile-reactive controllers (MPC and PD) and open-loop grasping. Our results show that LeTac-MPC has optimal performance in dynamic and force-interactive tasks and optimal generalizability. We release our code and dataset at https://github.com/ZhengtongXu/LeTac-MPC.
@@ I Introduction
Grasping is a reactive task in which humans estimate the states and physical properties of the grasped object by tactile feedback of the fingers and dynamically adjust finger behavior. This allows for robust grasping regardless of the object’s stiffness, whether it needs to be stationary or dynamically moved, and whether it is subjected to external forces. An ideal grasping algorithm for robots should be able to do the same. In this paper, we study the task of robotic tactile-reactive grasping. We specifically focus on a vision-based tactile sensor, GelSight [1], due to its advantages of perceiving high-resolution tactile feedback that contains rich information on the physical properties and states of the grasped object.
@@
1) Complex object physical properties: Objects may have different physical properties, which poses great challenges to tactile-based perception. Existing tactile perception methods often assume that grasped objects are rigid [10, 11], and sensors, such as vision-based tactile sensors [12, 13, 14, 15], may not be sensitive if the grasped object is softer than the sensor elastomer [1]. As a result, it can be difficult for controllers to get high-quality and stable feedback signals using model-based extraction methods when handling objects with different physical properties, and it can be challenging to generalize the same set of controller parameters to different objects, as their physical properties and dynamics may differ.
@@
2) Dynamic and force-interactive tasks: Statically grasping an object is often insufficient for dynamic and force-interactive tasks. For example, when robots use the gripper to pick fruit, there is a force interaction between the gripper, the fruit, and the branches. Similarly, in dynamic pick-and-place tasks, the grasped object may experience inertia force and collide with unexpected obstacles during transportation. In both of these scenarios, if a gripper is not responsive enough to state changes, external disturbances, and force interactions, it can result in the grasped object falling from the gripper. In addition, if a larger grasping force is used for a stronger grasp, delicate objects such as fruits, eggs, bread, and crackers can be destroyed. Therefore, it is challenging to design a reactive controller that can grasp with appropriate force according to the physical properties and states of the object.
@@
3) High-resolution tactile feedback integration: Another challenge is the integration of high-resolution tactile feedback, such as the vision-based tactile sensor GelSight [1], which contains rich information, into a real-time control loop for tactile-reactive grasping. Traditional control methods are based on low-dimensional feedback signals, which require the extraction of required information from tactile images for reduction of dimensionality [16, 17, 15]. However, many feature extraction methods make assumptions about the physical properties and shape of the object [10, 11, 16], making it difficult to generalize to objects with different physical properties and shapes. On the other hand, robot learning methods are based on high-dimensional observations such as visual and tactile images. However, these methods often overlook crucial aspects of the control process, such as convergence, response speed, control frequency, and constraints. As a result, they are more suitable for generating open-loop or low-frequency actions, but not well-suited for tasks that demand rapid and reactive responses.
@@
To address these challenges, we propose LeTac-MPC, a learning-based model predictive control (MPC) for tactile-reactive grasping. The novelties of the proposed approach are as follows.
@@
1) We design a neural network (NN) architecture with a differentiable MPC layer as the output layer. The MPC layer is modeled according to the control objectives to rationalize the embeddings extracted by the NN from the tactile feedback. The combination of NN and MPC layer provides the advantages of strong generalizability and ease of implementation as a real-time controller. We implement the trained NN and MPC layer in a model predictive controller that is convergent and can run at a frequency of 25 Hz. This controller enables robust reactive grasping of various daily objects in dynamic and force-interactive tasks.
@@
2) To train the model, we use four standardized blocks of different materials with different physical properties to collect a dataset and propose a fully automated data collection pipeline. Finally, we show that despite our data collection with standardized blocks in terms of sizes, shapes, and materials, and with no textures on the block surface, our trained controller can generalize to various daily objects with different sizes, shapes, materials, and textures. This shows the strong generalizability of LeTac-MPC.
@@
3) We also compare two purely model-based tactile-reactive controllers (MPC and PD) as baseline methods. Through experimental comparisons between these two model-based controllers, LeTac-MPC, and open-loop grasping, we show that LeTac-MPC has optimal performance in dynamic and force-interactive tasks and optimal generalizability.
@@ II-A Tactile Control
Previous research has shown that tactile feedback provided by vision-based tactile sensors can be used to design manipulation controllers for rigid objects [11, 18, 17], cables [16], and clothes [19]. These methods extract state estimations of the manipulated object, such as pose and contact line with the environment, from tactile feedback, and then design controllers. However, these approaches are not generalizable to objects with different physical properties due to limitations in their model-based state extraction methods. The work presented in [20] proposes a tactile control framework for robotic pushing that uses tactile feedback to estimate and predict state information. In [21], a learning-based method is proposed for flipping pages via tactile sensing. However, these two works do not fully use the rich information about the object’s physical properties that can be obtained from tactile feedback, which is critical for enhancing generalization to different objects.
Recently, the work in [22] proposes a method to learn soft tactile sensor membrane dynamics. However, this method does not consider the dynamics of the manipulated object, which is important for manipulating deformable objects. The work in [23] proposes a learning-based tactile MPC framework. However, this framework is not suitable for real-time control tasks because it uses a video prediction model, resulting in high computational cost. As mentioned in the paper, during the experimental testing, this MPC only runs at around 1 Hz.
@@ II-B Tactile Grasping
Multiple works exploit tactile feedback and NN to predict grasp success and stability [24, 25, 26, 27]. Based on the predictive ability of NN, the works in [28, 29, 30, 31] sample potential grasping actions and select the action with the lowest cost as the optimal grasping. However, these methods do not focus primarily on real-time control. They do not consider important aspects such as convergence, response speed, control frequency, and constraints in the control process. Consequently, they are better suited for low-frequency action generation rather than real-time control. For example, in [30], the computation times of different NN models for one trial are both more than 0.28 s. As a result, these methods cannot respond fast and smooth enough when the robot performs dynamic and force-interactive tasks. The works in [32, 10] maintain stable grasping by slip detection. Therefore, the accuracy of slip detection determines the performance of the method. However, in practical situations, slip detection may be insufficiently sensitive due to the different shapes and physical properties of grasped objects.
In [33, 16], grasping controllers with model-based tactile feature extraction are proposed. However, these methods cannot be widely applicable to objects with different shapes and physical properties since some feature extraction methods only work on specific types of object. For example, GelSight, as used in [33, 16], cannot obtain obvious features for objects softer than the sensor gel [1]. These features are extremely difficult to process and extract using model-based methods. Instead, existing work on tactile perception shows that NN can effectively extract these features [9]. Therefore, it is crucial to expand learning-based tactile perception to develop a learning-based tactile control method that can generalize to objects with different shapes, textures, and physical properties.
@@ II-C Learning Physical Properties for Manipulation
Learning-based methods that incorporate tactile and visual feedback can achieve dynamic manipulation without prior knowledge of the physical properties of the manipulated object [34, 35, 36]. However, these methods are not suitable for tasks that require reactive behavior. In contrast, many manipulation tasks require real-time sensing and real-time control. In [37, 38], the authors propose methods that utilize neural radiance fields (NeRF) for manipulating objects with complex properties. However, these methods are not suitable for tasks demanding real-time control and fast response due to the relatively high computational costs. In [39], a model learning method for elastic deformable linear objects is proposed that enables real-time cable manipulation. The input of its model is the cable state extracted from the raw image, which requires markers on the cable for tracking. However, in practice, extracting cable states from images without markers is challenging. In summary, achieving real-time, learning-based robot control directly using high-dimensional images as input remains a significant challenge.
@@ III Approach
This section outlines our proposed learning-based MPC for tactile-reactive grasping. Initially, we design an NN architecture that includes a differentiable MPC layer (Sections III-A and III-B) to extract and represent the diverse and complex physical properties of grasped objects. Then, we introduce the details of training our proposed model (Section III-C). Furthermore, we describe our automated data collection pipeline in Section III-D. Finally, we present the implementation of the NN as a real-time controller in Section III-F.
By our method, the gripper can successfully grasp objects with different physical properties and shapes while applying the appropriate force to avoid damage to fragile and delicate items. Our method can also adjust the grasping in real-time based on tactile feedback to improve the grasping robustness when performing dynamic and force-interactive tasks.
@@ III-A Tactile Information Encoding
We use GelSight [1], a vision-based tactile sensor that provides high-resolution images of the contact surface geometry and strain field, as shown in Fig. LABEL:fig:firstPage. To extract states and physical properties of grasped objects, we use a convolutional neural network (CNN) with a multi-layer perceptron (MLP) to encode a tactile image into a low-dimensional embedding $\mathbf{f}\in\mathbb{R}^{M}$, where $M$ is the dimension of the embedding. The works on tactile perception in [9, 41] demonstrate effective feature extraction from tactile images using pre-trained visual models. In our implementation, we use a pre-trained ResNet-152 [42] as the CNN architecture and replace the last layer with a two-layer MLP with ReLU activation.
@@ III-B Differentiable MPC Layer
In this section, we introduce a differentiable optimization layer that utilizes the extracted embeddings for controller design, as depicted in Fig. 1. This layer formulates an MPC problem, which we henceforth refer to as a differentiable MPC layer.
Differentiable optimization layers [43] function similarly to standard layers in neural networks, capable of performing both forward passes and backpropagation in batch form. During the forward pass, the layer takes in variables that define an optimization problem and outputs results related to the solution of that problem. Consequently, the forward pass of this layer involves solving optimization problems in batch form. Backpropagation in the differentiable optimization layer involves updating the parameters of the optimization problem based on the calculated gradients.
In this paper, we consider the two-finger gripper with one degree of freedom linear motion. This kind of gripper is widely used because of its practicality. Denoting the sampling time interval as $\Delta t$, we can first formulate the gripper motion model as:
@@
Scalars $p,v$, and $a$ are the position, velocity, and acceleration of the motion of gripper fingers. Note that $p,v$, and $a$ do not refer to the motion of the gripper itself in Cartesian space but rather denote the one-degree-of-freedom motion of the gripper fingers. The right subscript $n$ represents the $n$-th time step.
@@
where $\mathbf{A}_{f}\in\mathbb{R}^{M}$ would be learned. This modeling approach does not lose its applicability to complex dynamics because the tactile embeddings are actually generated by an NN, which has strong nonlinear fitting capabilities. Moreover, for MPC, we do not expect the model to maintain high accuracy over long timescales because MPC operates at a high frequency with receding horizon control. Therefore, as long as the model demonstrates adequate representational capability in a short-term local area, it is sufficient for MPC performance. The combination of the NN and the MPC layer provides the advantages of strong generalizability and ease of implementation as a real-time controller. The forward pass of the MPC layer involves solving the following optimization problem:
@@
$n$ denotes the current time step, and we can define the initial time step as $n=0$. As time progresses, there is no upper limit to $n$. Therefore, $n=0,1,2,3,\ldots$. $\mathbf{a}_{n}=[a_{n},a_{n+1},\ldots{},a_{n+N-1}]^{T}\in\mathbb{R}^{N}$ is the acceleration sequence, $N$ is the prediction horizon, the vector $\mathbf{A}_{f}\in\mathbb{R}^{M}$ is part of the state transition matrix, $\mathbf{Q}_{f}\in\mathbb{R}^{M\times M}$ is the weight matrix for $\mathbf{f}_{n}$, and $Q_{v}$ and $Q_{a}$ are the weight coefficients for $v_{n}$ and $a_{n}$, respectively. The scalar $P$ is used to amplify the terminal cost to speed up convergence. Note that the “$n$” in $v_{n},v_{n+1}$ represents the current time step $n$, while the “$N$” in $v_{n+N}$ represents the MPC horizon length $N$. Importantly, we design $\mathbf{f}_{n+1}$ is only directly related to $\mathbf{f}_{n}$ and $v_{n}$. This design helps the model avoid the influence of the different sizes of the grasped objects. Therefore, the trained model can be generalized to objects with different sizes naturally.
Upon solving the optimization problem (III-B) to obtain $\mathbf{a}_{n}^{*}=[a_{n}^{*},a_{n+1}^{*},\ldots,a_{n+N-1}^{*}]^{T}\in\mathbb{R}^{N}$, we can generate the entire trajectory of the gripper motion output by the MPC controller, including $p^{*},v^{*}$ and $a^{*}$, using equation (5). Ultimately, the final output of the MPC layer consists of the position sequence $[p^{*}_{n+1},\ldots,p^{*}_{n+N}]^{T}\in\mathbb{R}^{N}$, as shown in Fig. 1. The rationale behind this design is that most grippers, such as the WSG gripper used in this paper, have a low-level tracking controller capable of effectively tracking high-level position commands. By outputting the position sequence of the trajectory generated by the MPC layer and sending it to the low-level controller, we can approximate the tracking of the entire trajectory, including $p^{*},v^{*}$ and $a^{*}$, in a simple yet effective manner. For more details, see the experiments described in Section V.